HETAV PATEL

Grade 11,

Find A and B if the following equation if of a circle : (Ax)^2 + (B+4)y^2 + (A-3)xy + 5x + 7y + 5 = 0

Find A and B if the following equation if of a circle :
(Ax)^2 + (B+4)y^2 + (A-3)xy + 5x + 7y + 5 = 0

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
If the equation is of circle :
xy term should be zero : A-3 = 0 , Hence A = 3
The coefficient of x^2 and y^2 should be equal :
A^2 = B + 4
Hence, B = 5 , A = 3
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty